# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

### v2

@@ A.2 Cross-fitting and folds
Let $F:\{1,\dots,n\}\!\to\!\{1,\dots,K\}$ be a deterministic fold map (e.g., a hash of x_idx\_id). For any learner $\mathcal{L}$, train $\hat{\eta}^{(-k)}=\mathcal{L}$ on $\{i:F(i)\neq k\}$ and use out-of-fold predictions $\hat{\eta}^{\mathrm{OOF}}_{i}=\hat{\eta}^{(-F(i))}(O_{i})$ in influence-function (IF) calculations. The same folds are reused across AutoCal-R, SIMCal-W, and DR nuisances.
@@ A.3 Projection operators used by CJE
Simplex hull. For centered IF columns $\{\phi^{(e)}\}_{e\in\mathcal{E}}$, let $\Phi=[\phi^{(e)}]$ and $\Delta=\{\alpha:\alpha_{e}\geq 0,\ \sum_{e}\alpha_{e}=1\}$. IF-space stacking solves $\min_{\alpha\in\Delta}\alpha^{\top}\hat{\Sigma}\,\alpha$ with $\hat{\Sigma}=(1/n)\Phi^{\top}\Phi+\lambda I$.
A.4 AutoCal-R and SIMCal-W primitives
AutoCal-R. On $O=\{i:L_{i}=1\}$, fit $R=\hat{f}(Z(S))$ by either: (i) monotone ($Z(S)=S$), or (ii) two-stage ($Z(S,X)=\mathrm{ECDF}\{g(S,X)\}$ with a spline $g$), selecting by OOF RMSE (1-SE rule). Use OOF predictions $R^{\mathrm{OOF}}$ along IF paths; the point estimate may use a pooled fit. A terminal isotonic step enforces slice-mean preservation.
@@ A.4 AutoCal-R and SIMCal-W primitives
SIMCal-W. (Per fold) Fit up/down isotonic maps on $S$ to obtain $W^{\mathrm{OOF}}_{\uparrow},W^{\mathrm{OOF}}_{\downarrow}$; optionally include $W^{\mathrm{OOF}}_{\mathrm{base}}\equiv 1$. Define residuals $\Delta_{i}$ (IPS: $R_{i}$; DR: $R_{i}-\hat{q}(X_{i},A_{i})$). Choose $\hat{\beta}\in\arg\min_{\beta\in\Delta_{3}}\beta^{\top}\hat{\Sigma}\,\beta$ where $\hat{\Sigma}$ is the covariance of $U_{c}=W^{\mathrm{OOF}}_{c}\Delta$. Form $W^{\mathrm{stack}}=\sum_{c}\hat{\beta}_{c}W^{\mathrm{OOF}}_{c}$, renormalize to mean one (global, over the evaluation cohort), optionally apply the absolute variance guard
@@ A.6 Influence functions and variance
A.7 Oracle–uncertainty–aware (OUA) jackknife
Partition $O$ into $K$ oracle folds $\{O_{k}\}_{k=1}^{K}$. For each $k$, refit AutoCal-R on $O\setminus O_{k}$, recompute $R^{(-k)}$, and rerun the full pipeline to obtain $\hat{\psi}^{(-k)}$. Then
@@ A.9 Symbol glossary
$R=\hat{f}(Z(S))$ Calibrated reward (AutoCal-R; monotone or two-stage)
$\hat{W}_{\pi^{\prime}}$ Calibrated, unit-mean, $S$-monotone weights (SIMCal-W)
@@ Appendix B Algorithms (extended)
This appendix gives compact, cross-fitted pseudocode for CJE modules: reward calibration (AutoCal-R), surrogate-indexed weight calibration (SIMCal-W), estimators (Cal-IPS, DR-CPO), IF-space stacking, oracle–uncertainty–aware variance (OUA), and cluster bootstrap for Direct mode (Bootstrap-Direct). We reuse the same $K$-fold map $F(i)\!\in\!\{1{:}K\}$ across all modules. “OOF” denotes out-of-fold predictions used along the IF path.
Algorithm 1 AutoCal-R: mean-preserving reward calibration (cross-fitted; automatic two-stage fallback)
@@
8: Compute OOF risks (overall and by $S$ tertile); select mode via 1-SE rule (prefer monotone unless two-stage is significantly better)
@@
Algorithm 2 SIMCal-W: surrogate-indexed, unit-mean monotone weight calibration (OOF project $\to$ stack $\to$ cap $\to$ normalize)
@@
Algorithm 4 DR-CPO: sequence-aware doubly robust estimator (cross-fitted)
@@
Algorithm 5 IF-Stack: variance-optimal convex ensembling in IF space
@@
Algorithm 6 OUA jackknife: oracle–uncertainty–aware variance addition
@@
4:  Refit AutoCal-R on $O\setminus O_{k}$; recompute $R^{(-k)}$ and all downstream nuisances & weights; run the full pipeline to get $\widehat{V}^{(-k)}$
@@
PAVA is $O(n)$ after a shared sort by $S$. SIMCal-W is linear-time per fold; the stacking QP is $3{\times}3$ (weights) or a small $|\mathcal{E}|{\times}|\mathcal{E}|$ system. DR-CPO adds one rollout + judge per $(X,\pi^{\prime})$. The OUA jackknife refits the calibrator $K$ times and reruns the pipeline. Bootstrap-Direct requires B=2, [^0] B{=}2{,}000 refits of AutoCal-R; since isotonic regression is $O(n\log n)$, total cost is $O(Bn\log n)$. TF caches and precomputed features amortize cost.
@@ C.6 Proof of 1 (Cal-IPS)
where $\mathrm{rem}$ collects second-order sample-splitting terms. The empirical process term is $O_{p}(n^{-1/2})$; the second and third vanish by $L^{2}$-consistency of SIMCal-W and AutoCal-R (monotone or two-stage) and Cauchy–Schwarz; the remainder is $o_{p}(1)$ by cross-fitting. Finite-sample dispersion control follows from Lemma 2 and, if used, the blend cap $\rho\geq 1$; under the exact $\mathrm{IsoMeanOne}_{S}$ projection, $\mathrm{ESS}(\hat{W}_{\pi^{\prime}})\geq\mathrm{ESS}(W_{\pi^{\prime}}^{\mathrm{m1}})$.
C.7 Proof of Theorem 3 (DR-CPO $\sqrt{n}$ limits)
@@ C.8 Proof of Theorem 4 (budgeted bound)
Let $\mathcal{W}_{\rho}=\{m:\mathbb{E}[m]=1,\ m\uparrow S,\ \mathbb{E}[(m-1)^{2}]\leq\rho\,\mathbb{E}[(m^{\star}-1)^{2}]\}$. Intersecting the surrogate tangent space with the linear span induced by $m\in\mathcal{W}_{\rho}$ replaces $m^{\star}$ by its $L^{2}(P_{S})$ projection $m^{\star}_{\rho}=\Pi_{\mathcal{W}_{\rho}}(m^{\star})$ in $\phi_{\mathrm{sur}}$, giving $\phi^{(\rho)}$. Monotonicity in $\rho$ follows from nested convex sets $\mathcal{W}_{\rho_{1}}\subseteq\mathcal{W}_{\rho_{2}}$ for $\rho_{1}\leq\rho_{2}$, and $\lim_{\rho\to\infty}\phi^{(\rho)}=\phi_{\mathrm{sur}}$. If SIMCal-W converges to $m^{\star}_{\rho}$ (with the same cap), DR-CPO attains $\operatorname{Var}(\phi^{(\rho)})$ by the same one-of-two rate argument.
C.9 Proof of Theorem 5 (IF-space stacking) and Corollary 2
@@
C.10 Proof of 2 (OUA jackknife)
@@ C.11 Auxiliary lemmas used in the main proofs
Lemma 3 (OOF mean preservation for AutoCal-R).
Let $K$ be fixed and let $R^{\mathrm{OOF}}_{i}=\hat{f}^{(-F(i))}(Z(S_{i}))$ be OOF predictions from AutoCal-R (either mode). Then $P_{n}[R^{\mathrm{OOF}}]-P_{n}[Y]=o_{p}(1)$ and $P[R^{\mathrm{OOF}}-R^{\star}]=o_{p}(1)$ under $L^{2}(P_{S})$-consistency of $\hat{f}$.
@@
Let $\widehat{V}_{\mathrm{DR}}$ be DR-CPO with cross-fitted $(\hat{q},\hat{g}_{\pi^{\prime}})$ and calibrated $\hat{W}_{\pi^{\prime}}$. Then
@@
Metric projections onto closed convex sets in Hilbert spaces are firmly non-expansive: $\|\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y)\|^{2}\leq\langle\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y),\,x-y\rangle$. Hence compositions of DbP modules (reward, weight, IF-space projections) are non-expansive. See, e.g., Bauschke & Combettes (2017, Prop. 4.16).
@@
C.12 Coverage-Limited Efficiency (CLE) Diagnostic
We derive the CLE diagnostic stated in Equation 1. Let $\mathcal{T}\subset\mathcal{X}\times\mathcal{A}$ be a target-relevant region with $\alpha=P_{\pi^{\prime}}(\mathcal{T})$ and $\beta=P_{\pi_{0}}(\mathcal{T})$.
@@ C.12 Coverage-Limited Efficiency (CLE) Diagnostic
Proposition 4 (CLE diagnostic for IPS-style estimators).
@@
The CLE diagnostic has three multiplicative factors: (i) the coverage penalty $\alpha/\sqrt{\beta}$, which explodes when the logger rarely visits target-typical regions ($\beta\ll\alpha$); (ii) the shape mismatch $\sqrt{1+\chi^{2}}$, which inflates the floor even with good coverage if the target concentrates differently within $\mathcal{T}$; (iii) the standard noise term $\sigma_{\mathcal{T}}/\sqrt{n}$.
@@
C.13 EIF derivation for Direct mode
@@
C.14 Assumptions ledger
For any CJE deployment, record and validate the following assumptions:
@@ C.14 Assumptions ledger
$\mathbb{E}[Y^{*}\mid X,A]=\mathbb{E}[Y\mid X,A]$. The operational oracle $Y$ is sufficiently aligned with the idealized deliberation oracle $Y^{*}$ such that optimizing for $Y$ approximates optimizing for $Y^{*}$.
Validation: Bridge Validation Protocol (BVP) with predictive transportability experiments, construct validity audits, and stability monitoring.
S1 (Prentice Sufficiency).
$\mathbb{E}[Y\mid X,A,S]=f(S,X)$ a.s. on $\operatorname{supp}(\pi_{0}\cup\Pi_{\mathrm{eval}})$. The surrogate $S$ captures all relevant information about the oracle $Y$ given covariates $X$.
Testable implication: Residuals $Y-\hat{f}(S,X)$ should be mean-zero and uncorrelated with $A$ conditional on $(X,S)$.
S2 (Transportability).
The calibration function $f$ transports across environments $g\in\mathcal{G}$: $\mathbb{E}_{g}[Y\mid X,A,S]=f(S,X)$ for all $g$ (Pearl & Bareinboim, 2014).
Graphical test: S-admissibility requires that no selection node points into $Y$ given $S$ in the selection diagram with incoming arrows to $A$ removed.
Testable implication: Calibration curves should be stable across (i) policy subgroups, (ii) time periods, and (iii) covariate strata.
S3 (Overlap/Positivity).
$\pi^{\prime}(a\mid x)>0\Rightarrow\pi_{0}(a\mid x)>0$ a.s. Required for off-policy estimation.
@@
Diagnostic: Target-Typicality Coverage (TTC). If TTC <0.7<0.7, expect logs-only IPS to fail.
L1 (Oracle MAR).
$L\perp Y\mid(X,A,S)$, where $L\in\{0,1\}$ indicates oracle labeling. Oracle labeling is ignorable conditional on observed surrogates.
Testable implication: Oracle-labeled and unlabeled samples should have similar $(X,A,S)$ distributions.
L2 (Oracle Positivity).
$P(L=1\mid X,A,S)>0$ on the support where calibration will be applied.
Diagnostic: Check for gaps in oracle labeling across the $(X,A,S)$ space.
C.15 What the bounds do not include

### v3

@@ A.2 Cross-fitting and folds
Let $F:\{1,\dots,n\}\!\to\!\{1,\dots,K\}$ be a deterministic fold map (e.g., a hash of x_idx\_id). For any learner $\mathcal{L}$, train $\hat{\eta}^{(-k)}=\mathcal{L}$ on $\{i:F(i)\neq k\}$ and use out-of-fold predictions $\hat{\eta}^{\mathrm{OOF}}_{i}=\hat{\eta}^{(-F(i))}(O_{i})$ in influence-function (IF) calculations. The same folds are reused across reward calibration, weight stabilization, and DR nuisances.
@@ A.3 Projection operators used by CJE
Simplex hull. For centered IF columns $\{\phi^{(e)}\}_{e\in\mathcal{E}}$, let $\Phi=[\phi^{(e)}]$ and $\Delta=\{\alpha:\alpha_{e}\geq 0,\ \sum_{e}\alpha_{e}=1\}$. Influence-function stacking solves $\min_{\alpha\in\Delta}\alpha^{\top}\hat{\Sigma}\,\alpha$ with $\hat{\Sigma}=(1/n)\Phi^{\top}\Phi+\lambda I$.
A.4 Reward calibration and weight stabilization primitives
Reward calibration. On $O=\{i:L_{i}=1\}$, fit $R=\hat{f}(Z(S))$ via: (i) two-stage (recommended): $Z(S,X)=\mathrm{ECDF}\{g(S,X)\}$ with a spline $g$, or (ii) monotone-only: $Z(S)=S$ when covariates are unavailable. Use OOF predictions $R^{\mathrm{OOF}}$ along IF paths; the point estimate may use a pooled fit. A terminal isotonic step enforces slice-mean preservation.
@@ A.4 Reward calibration and weight stabilization primitives
Weight stabilization. (Per fold) Fit up/down isotonic maps on $S$ to obtain $W^{\mathrm{OOF}}_{\uparrow},W^{\mathrm{OOF}}_{\downarrow}$; optionally include $W^{\mathrm{OOF}}_{\mathrm{base}}\equiv 1$. Define residuals $\Delta_{i}$ (IPS: $R_{i}$; DR: $R_{i}-\hat{q}(X_{i},A_{i})$). Choose $\hat{\beta}\in\arg\min_{\beta\in\Delta_{3}}\beta^{\top}\hat{\Sigma}\,\beta$ where $\hat{\Sigma}$ is the covariance of $U_{c}=W^{\mathrm{OOF}}_{c}\Delta$. Form $W^{\mathrm{stack}}=\sum_{c}\hat{\beta}_{c}W^{\mathrm{OOF}}_{c}$, renormalize to mean one (global, over the evaluation cohort), optionally apply the absolute variance guard
@@ A.6 Influence functions and variance
A.7 Calibration-aware jackknife
Partition $O$ into $K$ oracle folds $\{O_{k}\}_{k=1}^{K}$. For each $k$, refit reward calibration on $O\setminus O_{k}$, recompute $R^{(-k)}$, and rerun the full pipeline to obtain $\hat{\psi}^{(-k)}$. Then
@@ A.9 Symbol glossary
$R=\hat{f}(Z(S))$ Calibrated reward (monotone or two-stage)
$\hat{W}_{\pi^{\prime}}$ Stabilized, unit-mean, $S$-monotone weights
@@ Appendix B Algorithms (extended)
This appendix gives compact, cross-fitted pseudocode for CJE modules: reward calibration, weight stabilization, estimators (calibrated IPS, calibrated DR), influence-function stacking, calibration-aware variance estimation, and cluster bootstrap for Direct mode. We reuse the same $K$-fold map $F(i)\!\in\!\{1{:}K\}$ across all modules. “OOF” denotes out-of-fold predictions used along the IF path.
Algorithm 1 Reward-Calibration: mean-preserving calibration (cross-fitted; automatic two-stage fallback)
@@
8: Default: use two-stage mode (recommended when covariates like response length are available); use monotone-only as fallback
@@
Algorithm 2 Weight-Stabilization: surrogate-indexed, unit-mean monotone calibration (OOF project $\to$ stack $\to$ cap $\to$ normalize)
@@
Algorithm 4 Calibrated-DR: sequence-aware doubly robust estimator (cross-fitted)
@@
Algorithm 5 Influence-Function Stacking: variance-optimal convex ensembling in influence-function space
@@
Algorithm 6 Cal-Aware-Jackknife: calibration-uncertainty-aware variance addition
@@
4:  Refit reward calibration on $O\setminus O_{k}$; recompute $R^{(-k)}$ and all downstream nuisances & weights; run the full pipeline to get $\widehat{V}^{(-k)}$
@@
PAVA is $O(n)$ after a shared sort by $S$. Weight stabilization is linear-time per fold; the stacking QP is $3{\times}3$ (weights) or a small $|\mathcal{E}|{\times}|\mathcal{E}|$ system. Calibrated DR adds one rollout + judge per $(X,\pi^{\prime})$. The calibration-aware jackknife refits the calibrator $K$ times and reruns the pipeline. Bootstrap-Direct requires B=2, [^0] B{=}2{,}000 refits of the reward calibrator; since isotonic regression is $O(n\log n)$, total cost is $O(Bn\log n)$. TF caches and precomputed features amortize cost.
@@ C.6 Proof of 2 (Cal-IPS)
where $\mathrm{rem}$ collects second-order sample-splitting terms. The empirical process term is $O_{p}(n^{-1/2})$; the second and third vanish by $L^{2}$-consistency of weight stabilization and reward calibration (monotone or two-stage) and Cauchy–Schwarz; the remainder is $o_{p}(1)$ by cross-fitting. Finite-sample dispersion control follows from Lemma 2 and, if used, the blend cap $\rho\geq 1$; under the exact $\mathrm{IsoMeanOne}_{S}$ projection, $\mathrm{ESS}(\hat{W}_{\pi^{\prime}})\geq\mathrm{ESS}(W_{\pi^{\prime}}^{\mathrm{m1}})$.
C.7 Proof of Theorem 3 (Calibrated DR $\sqrt{n}$ limits)
@@ C.8 Proof of Theorem 4 (budgeted bound)
Let $\mathcal{W}_{\rho}=\{m:\mathbb{E}[m]=1,\ m\uparrow S,\ \mathbb{E}[(m-1)^{2}]\leq\rho\,\mathbb{E}[(m^{\star}-1)^{2}]\}$. Intersecting the surrogate tangent space with the linear span induced by $m\in\mathcal{W}_{\rho}$ replaces $m^{\star}$ by its $L^{2}(P_{S})$ projection $m^{\star}_{\rho}=\Pi_{\mathcal{W}_{\rho}}(m^{\star})$ in $\phi_{\mathrm{sur}}$, giving $\phi^{(\rho)}$. Monotonicity in $\rho$ follows from nested convex sets $\mathcal{W}_{\rho_{1}}\subseteq\mathcal{W}_{\rho_{2}}$ for $\rho_{1}\leq\rho_{2}$, and $\lim_{\rho\to\infty}\phi^{(\rho)}=\phi_{\mathrm{sur}}$. If weight stabilization converges to $m^{\star}_{\rho}$ (with the same cap), calibrated DR attains $\operatorname{Var}(\phi^{(\rho)})$ by the same one-of-two rate argument.
C.9 Proof of Theorem 5 (influence-function stacking) and Corollary 2
@@
C.10 Proof of 3 (calibration-aware jackknife)
@@ C.11 Auxiliary lemmas used in the main proofs
Lemma 3 (OOF mean preservation for reward calibration).
Let $K$ be fixed and let $R^{\mathrm{OOF}}_{i}=\hat{f}^{(-F(i))}(Z(S_{i}))$ be OOF predictions from reward calibration (either mode). Then $P_{n}[R^{\mathrm{OOF}}]-P_{n}[Y]=o_{p}(1)$ and $P[R^{\mathrm{OOF}}-R^{\star}]=o_{p}(1)$ under $L^{2}(P_{S})$-consistency of $\hat{f}$.
@@
Let $\widehat{V}_{\mathrm{DR}}$ be calibrated DR with cross-fitted $(\hat{q},\hat{g}_{\pi^{\prime}})$ and calibrated $\hat{W}_{\pi^{\prime}}$. Then
@@
Metric projections onto closed convex sets in Hilbert spaces are firmly non-expansive: $\|\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y)\|^{2}\leq\langle\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y),\,x-y\rangle$. Hence compositions of projection modules (reward, weight, IF-space projections) are non-expansive. See, e.g., Bauschke & Combettes (2017, Prop. 4.16).
@@
C.12 Coverage-Limited Efficiency (CLE) bound
We derive the CLE bound stated in Equation 1. Let $\mathcal{T}\subset\mathcal{X}\times\mathcal{A}$ be a target-relevant region with $\alpha=P_{\pi^{\prime}}(\mathcal{T})$ and $\beta=P_{\pi_{0}}(\mathcal{T})$.
@@ C.12 Coverage-Limited Efficiency (CLE) bound
Proposition 5 (CLE bound for IPS-style estimators).
@@
The CLE bound has three multiplicative factors: (i) the coverage penalty $\alpha/\sqrt{\beta}$, which explodes when the logger rarely visits target-typical regions ($\beta\ll\alpha$); (ii) the shape mismatch $\sqrt{1+\chi^{2}}$, which inflates the floor even with good coverage if the target concentrates differently within $\mathcal{T}$; (iii) the standard noise term $\sigma_{\mathcal{T}}/\sqrt{n}$.
@@
C.13 TTC collapse with sequence length (structural coverage sparsity)
Even with perfect propensity estimates (no teacher-forcing noise), logger coverage collapses exponentially with sequence length. This formalizes why logs-only OPE is fundamentally impractical for open-ended generation, independent of TF brittleness.
Proposition 6 (Exponential collapse of logger coverage).
Let $q$ and $p$ be token-level distributions with $\mathrm{KL}(p\|q)=\delta>0$ per token. For i.i.d. sequences of length $L$, the trajectory-level KL divergence is $L\delta$, and
$\mathrm{TTC}:=P_{\pi_{0}}(\text{target-typical region})\;\leq\;\exp\bigl(-\Omega(L\delta)\bigr).$
When $L\delta\gg 1$, logger coverage is exponentially small in $L$.
Sketch.
Under the product measure, trajectory log-likelihood ratios concentrate around $L\cdot\mathbb{E}_{p}[\log(p/q)]=L\delta$. By standard large-deviation arguments (Sanov’s theorem), the probability that a logger-sampled trajectory falls in the target-typical region decays as $\exp(-L\cdot D)$ for some $D>0$ depending on $\delta$. ∎
Simulation verification.
[…]
Implication. For open-ended LLM generation with typical response lengths ($L\approx 100$$500$ tokens), even tiny per-token KL divergence ($\delta\approx 0.03$ nats) yields negligible TTC. This is not a TF bug—it is the geometry of high-dimensional trajectory spaces. Logs-only OPE faces a structural precision floor that no estimator can overcome.
C.14 EIF derivation for Direct mode
@@
C.15 Assumptions and what CJE relaxes
CJE’s key innovation is replacing untestable identification assumptions with auditable diagnostics. We organize assumptions into four categories, with mean transport as the fundamental identification requirement.
@@ C.15 Assumptions and what CJE relaxes
$\mathbb{E}[Y^{*}\mid X,A]=\mathbb{E}[Y\mid X,A]$. The operational oracle $Y$ is sufficiently aligned with the idealized deliberation oracle $Y^{*}$ such that optimizing for $Y$ approximates optimizing for $Y^{*}$. This is a governance question outside the statistical framework—oracle selection requires domain expertise and construct validity audits.
Category A: Design/missingness requirements (always needed).
  • Oracle MAR (L1): $L\perp Y\mid(X,A,S)$. Oracle labeling is ignorable conditional on observed data.
  • Oracle positivity (L2): $P(L=1\mid X,A,S)>0$ on the support where calibration will be applied.
Category B: Identification for surrogate-only mean claims (the fundamental requirement).
The requirement for valid policy value estimation is mean transport:
$\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0.$
This is exactly the condition tested by the mean residual audit (4). CJE treats transport as testable rather than assumed:
  • With audit (${\sim}50$$200$ oracle labels per $\pi^{\prime}$): Test transport directly. Pass $\to$ reuse calibration; fail $\to$ recalibrate or refuse level claims.
  • Without audit: Transport must be assumed. In our experiments, calibration transported well across comparable policies (clone, premium, prompt-engineered) but failed for adversarial ones. How aggressively you audit depends on context—stakes, policy similarity, and historical transport stability.
Category C: Structural assumptions (sufficient for transport, not required).
The following are sufficient conditions that imply transport. CJE does not require them when transport is verified via audit:
  • Prentice-style surrogacy: $Y\perp A\mid(X,S)$ (distributional stability across policies).
  • Mean sufficiency / single-index (J2-M, J2-SI): $\mathbb{E}[Y\mid X,A,S]=\mu(g(S,X))$ for some monotone $\mu$ and index $g$.
These structural assumptions also enable the efficiency theory (Theorems 1 and 3), but are not required for consistent estimation when transport is audited. Mean sufficiency implies transport (if it holds globally), but transport can hold without it—errors may cancel in expectation even when the structural assumption fails.
Category D: Estimator-dependent requirements.
Overlap (S3): $\pi^{\prime}(a\mid x)>0\Rightarrow\pi_{0}(a\mid x)>0$. Required for IPS and DR estimators. The Direct method requires no overlap—it generates fresh responses under $\pi^{\prime}$.
@@
Diagnostic: Target-Typicality Coverage (TTC). If TTC <0.7<0.7, logs-only IPS will fail due to coverage-limited efficiency; prefer Direct.
C.16 Additional corollaries
Corollary 1 (Blackwell-efficiency monotonicity).
If $S_{2}$ is a garbling of $S_{1}$ (i.e., $\sigma(S_{2})\subseteq\sigma(S_{1})$), then $\operatorname{Var}\!\big(\phi_{\mathrm{sur}}(S_{1})\big)\leq\operatorname{Var}\!\big(\phi_{\mathrm{sur}}(S_{2})\big)$, with strict inequality unless $W_{\pi^{\prime}}$ is already $\sigma(S_{2})$-measurable and $R^{\star}$ is degenerate. This is a statistical analog of Blackwell’s theorem on experiment comparisons (Blackwell, 1953): finer information structures dominate coarser ones.
Corollary 2 (Carathéodory sparsity).
C.17 What the bounds do not include